OpenStudy (anonymous):

solve for x 9*x = square root 729

OpenStudy (lgbasallote):

maybe it's written as \[9^x = \sqrt{729}?\]

OpenStudy (anonymous):

yes lgbasallote you right

OpenStudy (lgbasallote):

lol then 729 = 9^3 sqrt 729 = 3/2

OpenStudy (lgbasallote):

9^x = 9^(3/2)

OpenStudy (lgbasallote):

you get the logic?

OpenStudy (lgbasallote):

note that: \[a^x = a^y \implies x = y\]

OpenStudy (anonymous):

9^x = √729 9^x = 27 3^2x = 3^3 Now just solve for x in the exponents since the bases are the same. 2x = 3 x = 3/2 or 1.5